A cruise ship with a mass of 1.1 × 107 kg strikes a pier at a speed of 0.65 m/s. It comes to rest over a distance of 5.9 m, damaging the ship, the pier, and the tugboat captain’s finances. Assume a constant acceleration occurs during the collision.

Find the average force Fsp being exerted on the pier as the ship hits it, in terms of the time of interaction Δt, the initial velocity of the ship v, and the mass of the ship m. Take the motion of the cruise ship to be in the positive direction.
Answer #1

  in terms of the time of interaction Δt, the initial velocity of the ship v, and the mass of the ship m

Fsp = m*Δv /Δt  

t = d/v1 = 5.9 /0.65 = 9.07692308 sec

the force is retained so total time = 2*t = 2*9.07692308 =18.1538462 s  

F = 1.1*10^7 * ( vf-vi ) /Δt =   1.1*10^7 *( 0 - 0.65 ) / 18.1538462

F = 3.9385*105 N

Goodluck for exams Comment in case any doubt, will reply for sure..

we multiplied time by 2 because we have average velocity

x = x0 +( v-v0) /2 *t